OpenStudy (anonymous):

Simplify 7x^-8 * 6x^3

OpenStudy (anonymous):

\[6x^3*7x^-8=\frac{ 6x^3 }{ 7x^8 }=\frac{ 6 }{ 7x^5 }\]